# SIC-POVMs and orders of real quadratic fields

### v1

@@ 1.3. Main results
In this paper, we refine 1.1 to conjectures that specify the number fields associated to all the Weyl–Heisenberg SICs in each dimension $d\geq 4$ and which count the number of such SICs associated to each field.
More precisely, we connect both SICs and fields to certain orders of the real quadratic number field $K_{d}=\mathbb{Q}(\sqrt{\Delta_{d}})$, with $\Delta_{d}=(d+1)(d-3)$. Write $\Delta_{d}=f^{2}\Delta_{0}$, where $\Delta_{0}$ is the fundamental discriminant $K_{d}$, and $f\geq 1$. The orders that arise are exactly those $\mathcal{O}^{\prime}$ satisfying $\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{K}$ for $\varepsilon_{d}=\frac{1}{2}(d-1+\sqrt{\Delta_{d}})$. These orders, arranged by set inclusion, form a distributive lattice, isomorphic to that of the positive divisors of $f$ under divisibility.
We organize Weyl–Heisenberg SICs into equivalence classes. For a line-SIC $S$,
@@
    the geometric equivalence class $[S]$ is the orbit of $S$ under the action of the projective extended Clifford group $\operatorname{PEC}(d)$ (see 2.9), and
@@
    the Galois multiplet $[[S]]$ is the set of geometric equivalence classes that are permuted with each other under a certain Galois action (see 4.11).
In Section 3, we define several fields attached to a line-SIC $S=\{\mathbb{C}\mathbf{v}_{k}:1\leq k\leq d^{2}\}$. These fields are finitely generated extensions of the rational numbers $\mathbb{Q}$.
@@
    The ratio SIC-field $F^{\rm vec}\!\left(S\right)$ (called the “SIC field” in [10]) is the field generated by all the ratios $\frac{v_{ki}}{v_{kj}}$ of distinct entries ($i\neq j$, $v_{kj}\neq 0$) of the vectors $\mathbf{v}_{k}$ with $\mathbf{v}_{k}\in S$.
@@
    The projection SIC-field $F^{\rm proj}\!\left(S\right)$ is the field generated by all the matrix entires of the Hermitian projection operators $\Pi_{i}=\mathbf{v}_{i}\mathbf{v}_{i}^{\dagger}$ of the $d^{2}$ unit vectors $\{{\mathbf{v}}_{i}:1\leq i\leq d^{2}\}$ defining $S$. It is shown to be the compositum of $F^{\rm vec}\!\left(S\right)$ with its complex conjugate field $\overline{F^{\rm vec}\!\left(S\right)}$ (in 3.3).
@@
    The extended projection SIC-field $F^{\rm eproj}\!\left(S\right)=F^{\rm proj}\!\left(S\right)\!(-e^{\frac{\pi i}{d}})$ is obtained by adjoining the root of unity $-e^{\frac{\pi i}{d}}$. For Weyl–Heisenberg SICs, we show $F^{\rm eproj}\!\left(S\right)$ is an invariant of the geometric equivalence class $[S]$ (in 3.8), and we show the index $[F^{\rm eproj}\!\left(S\right):F^{\rm proj}\!\left(S\right)]\in\{1,2\}$ and is $1$ if $d$ is odd (in 3.6).
@@
    For SICs defined over a finite extension of $\mathbb{Q}$, we define the Galois projection SIC-field $F^{\rm Gproj}\!\left(S\right)$ to be the normal closure of $F^{\rm eproj}\!\left(S\right)$, which is an invariant of the multiplet $[[S]]$. We show that the subgroup of $\operatorname{Gal}(F^{\rm Gproj}\!\left(S\right)/\mathbb{Q})$ consisting of elements that commute with complex conjugation sends SICs to other SICs with the same Galois projection SIC-field (in 4.8).
We have the field inclusions
@@
In 4.9, we show that 1.1 implies that
@@
The refinements of 1.1 formulated in this paper posit that Weil–Heisenberg SICs in dimension $d\neq 3$ occur in multiplets indexed by the intermediate orders $\mathcal{O}^{\prime}$ of $K_{d}$ the lie between $\mathcal{O}=\mathbb{Z}[\varepsilon_{d}]$ and $\mathcal{O}_{K_{d}}$. Each of these orders is in turn indexed by its conductor $f^{\prime}$, a positive integer such that $\operatorname{disc}(\mathcal{O}^{\prime})=(f^{\prime})^{2}\Delta_{0}$ where $\Delta_{0}=\operatorname{disc}(\mathcal{O}_{K_{d}})$. The conductors appearing are the positive divisors of the conductor $f$ of $\mathcal{O}$.
@@
Fix a positive integer $d\neq 3$, let $\Delta=\Delta_{d}=(d+1)(d-3)$, and write $\Delta_{d}=f^{2}\Delta_{0}$ for $\Delta_{0}=\operatorname{disc}(\mathcal{O}_{K_{d}})$ and some positive integer $f$. Then there is a bijection $\mathcal{M}$ between conductors of intermediate orders and Galois multiplets,
@@
    The number of geometric equivalence classes $[S]$ in the multiplet $\mathcal{M}(f^{\prime})=[[S]]$ is $\left|\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})\right|$, the class number of the invertible class group of the order $\mathcal{O}^{\prime}$ of discriminant $(f^{\prime})^{2}\Delta_{0}$.
@@
    If $f_{1}\mid f_{2}\mid f$, $[S_{1}]\in\mathcal{M}(f_{1})$, and $[S_{2}]\in\mathcal{M}(f_{2})$, then $F^{\rm vec}\!\left(S_{1}\right)\subseteq F^{\rm vec}\!\left(S_{2}\right)$.
A mildly strengthened version of this conjecture is stated as 7.2. The field inclusion property (2) puts a strong restriction on the map $\mathcal{M}$. Modulo the next conjecture, it determines $\mathcal{M}$ entirely for a fraction $\frac{47}{48}$ of all dimensions $d$ (in the sense of 9.10).
The following conjecture is an extension of 1.1 using the notion of ray class fields of an order $\mathcal{O}^{\prime}$ introduced in [33] and reviewed in Section 8.1.
@@
Let $d\geq 4$. Then for any Weyl–Heisenberg line-SIC $S$ in dimension $d$ with associated order $\mathcal{O}^{\prime}$ having discriminant $(f^{\prime})^{2}\Delta_{0}$ (that is, with $\mathcal{M}(f^{\prime})=[[S]]$),
$F^{\rm vec}\!\left(S\right)=H_{d^{\prime}\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}^{\prime}}$ (1.6)
@@
The special case $f^{\prime}=1$ corresponds to 1.1, where the order $\mathcal{O}^{\prime}$ is the maximal order. The field inclusions asserted in 1.1(2) follow from 1.2. Appleby, Flammia, McConnell, and Yard [9] call the associated Galois multiplet $[[S]]$ the minimal multiplet, as it has the smallest associated field $F^{\rm vec}\!\left(S\right)$ under inclusion. 8.4 gives a further refinement of 1.3; these conjectures refine and revise [31, Conj. V.17].
The paper describes numerical checks of predictions of these conjectures against numerical data. The data for $4\leq d\leq 90$ exhibits patterns consistent with grouping SICs into multiplets indexed by orders. We formulate several narrower conjectures that make numerical predictions that can be checked against existing data and give tables of comparisons.
We also derive some consequences of the refined conjectures. In Section 6, we show they imply that the number of distinct geometric equivalence classes $[S]$ will be larger than $1$ for all $d\geq 23$ and will grow like $d^{1+o(1)}$ as $d\to\infty$. (See 6.3 and 6.4.)
@@ 1.4. Notation
  • Boldface roman lower case letters represent column vectors in $\mathbb{C}^{d}$. For a complex vector, $\mathbf{w}^{\dagger}:=\overline{\mathbf{w}}^{T}$, and similarly for matrices.
@@ 2. Weyl–Heisenberg SICs
In this section, we defined and give background on Weyl–Heisenberg (covariant) SICs, which satisfy a particular group symmetry property.
@@ 2.1. Weyl–Heisenberg SICs
Specifically, Weyl–Heisenberg line-SICs will be defined as the orbit of a single line under the action of the Weyl-Heisenberg group, a finite subgroup of the group of $d\times d$ unitary matrices.
@@
The projective Weyl–Heisenberg group $\operatorname{PWH}(d)$ is defined to be $\operatorname{WH}(d)$ modulo its scalar matrix subgroup $Z(\operatorname{WH}(d))=\{\xi^{j}I:\,j\geq 0\}$, where $I$ is the $d\times d$ identity matrix.
@@
Its associated set of complex lines, denoted $\operatorname{WHL}(\mathbf{w})$ is
@@
We call a generating vector $\mathbf{v}$ of a Weyl–Heisenberg SIC a fiducial vector. Such a vector is well-defined up to multiplication by a complex scalar $c\in\mathbb{C}^{\times}$. Each SIC has a choice of $d^{2}$ fiducial vectors, up to scalars. We call two Weyl–Heisenberg SICs scalar-equivalent if they have a common fiducial vector, up to multiplication by a nonzero scalar.
@@
This conjecture is currently known, by explicit constructions to hold for finitely many dimensions $d$. At present it includes dimensions $1\leq d\leq 53$ and a extensive number of higher dimensions.
@@
For each dimension $d\geq 1$, except $d=3$, there are finitely many scalar-equivalence classes of Weyl–Heisenberg SICs.
There are a few small dimensions where this conjecture has been proved rigorously, including $d\in\{1,2,4,5\}$. In higher dimensions, up to at least $d=90$, it is supported by non-rigorous evidence given by numerical optimization routines repeatedly returning the same SICs [39, 38]. In dimension $3$, there is a one-parameter family of inequivalent Weyl–Heisenberg SICs giving uncountably many distinct scalar-equivalence classes of SICs.
A closely related conjecture is stated in Waldron [46, Sec. 14.27]: The (real) algebraic variety of (normalized Weyl–Heisenberg) SIC fiducials is zero-dimensional, except for $d=3$, where it is one-dimensional. This conjecture implies 2.5; it is equivalent if one takes “dimension” to mean “real dimension” (i.e., dimension of the Zariski closure) and is a stronger conjecture if one takes “dimension” to mean “dimension of the complex variety defined by the same equations.” Resolving 2.5 or Waldron’s conjecture in general (and specifically even for $n=6$) is an interesting open problem in optimization and real algebraic geometry.
@@ 2.2. Geometric equivalence classes of Weyl–Heisenberg SICs
There are two types of geometric operators acting on $\mathbb{C}^{d}$ that send SICs to SICs. Firstly, if $U$ is a unitary matrix in $\mathbf{U}(d)=\{U\in\mathbf{GL}(\mathbb{C}^{d}):UU^{\dagger}=1\}$ and $S=\{\mathbb{C}\mathbf{v}_{1},\ldots,\mathbb{C}\mathbf{v}_{d^{2}}\}$ is a SIC, then so is $US=\{\mathbb{C}U\mathbf{v}_{1},\ldots,\mathbb{C}U\mathbf{v}_{d^{2}}\}$. (This is a unitary equivalence.) Secondly, if $C_{d}$ is the complex conjugation operator acting pointwise on $\mathbb{C}^{d}$, so that $C_{d}\mathbf{v}:=\overline{\mathbf{v}}$, then $C_{d}$ also preserves the SIC property, as does any “antiunitary” operator of the form $C_{d}U$. These operators may be collected together to form the extended unitary group.
@@
The extended unitary group is $\mathbf{EU}(d):=\mathbf{U}(d)\sqcup C_{d}\mathbf{U}(d)$. The projective extended unitary group $\mathbf{PEU}(d)$ is defined to be $\mathbf{EU}(d)$ modulo its scalar matrix subgroup $\{zI:z\in\mathbb{C},\left|z\right|=1\}$.
@@
If $\mathbf{v}$ is a fiducial vector for a Weyl–Heisenberg covariant SIC and $U\in\mathbf{EU}(d)$, then the vector $U\mathbf{v}$ will be a fiducial vector for a SIC covariant for the group $G^{\prime}=U\operatorname{WH}(d)U^{-1}$. All currently known SICs are unitary-equivalent to a Weyl–Heisenberg group $\operatorname{WH}(d)$ covariant SIC, with a single exception, the Hoggar lines in $\mathbb{C}^{8}$ [29, 30]. The Hoggar lines are group covariant for the tensor product $\operatorname{WH}(2)\otimes\operatorname{WH}(2)\otimes\operatorname{WH}(2)$. (The group $\operatorname{WH}(2)\otimes\operatorname{WH}(2)\otimes\operatorname{WH}(2)$ may itself be thought of as a generalized (Weyl)-Heisenberg group.)
@@
The Clifford group $\operatorname{C}(d)$ is infinite because it contains a copy of $\mathbf{U}(1)$ as scalar matrices $zI$ with $\left|z\right|=1$; thus, $\operatorname{EC}(d)$ is infinite. However, $\operatorname{PC}(d)$ and $\operatorname{PEC}(d)$ are finite groups. (The extended Clifford $\operatorname{EC}(d)$ group is studied in connection with SICs by Appleby [1].)

### v3

@@ 1.3. Main results
In this paper, we refine 1.1 to conjectures that specify the number fields associated to all the Weyl–Heisenberg SICs in each dimension $d\geq 4$, and which also count the number of geometric equivalence classes of such SICs associated to each such field.
More precisely, we connect both SICs and their associated fields to certain orders of the real quadratic number field $K_{d}=\mathbb{Q}(\sqrt{\Delta_{d}})$, with $\Delta_{d}=(d+1)(d-3)$. Write $\Delta_{d}=f^{2}\Delta_{0}$, where $\Delta_{0}$ is the fundamental discriminant $K_{d}$, and $f\geq 1$. The orders that arise are exactly those $\mathcal{O}^{\prime}$ satisfying $\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{K}$ for $\varepsilon_{d}=\frac{1}{2}(d-1+\sqrt{\Delta_{d}})$. These orders, arranged by set inclusion, form a distributive lattice, isomorphic to that of the positive divisors of $f$ under divisibility.
We organize Weyl–Heisenberg SICs into equivalence classes. For such a line-SIC $S$,
@@
    the geometric equivalence class $[S]$ is the orbit of $S$ under the action of the projective extended Clifford group $\operatorname{PEC}(d)$ (by 2.9), and
@@
    the Galois multiplet $[[S]]$ is the set of geometric equivalence classes that are permuted with each other under a certain Galois action (by 4.9).
In Sections 3 and 4, we define several fields attached to a line-SIC $S=\{\mathbb{C}\mathbf{v}_{k}:1\leq k\leq d^{2}\}$. These fields are finitely generated extensions of the rational numbers $\mathbb{Q}$.
@@
    The ratio SIC field $F^{\rm vec}\!\left(S\right)$ (called the “SIC field” in [9]) is the field generated by all the ratios $\frac{v_{ki}}{v_{kj}}$ of distinct entries ($i\neq j$, $v_{kj}\neq 0$) of the vectors $\mathbf{v}_{k}$ with $\mathbf{v}_{k}\in S$.
@@
    The projection SIC field $F^{\rm proj}\!\left(S\right)$ is the field generated by the matrix entries of the Hermitian projection operators $\Pi_{i}=\mathbf{v}_{i}\mathbf{v}_{i}^{\dagger}$ of the $d^{2}$ unit vectors $\{{\mathbf{v}}_{i}:1\leq i\leq d^{2}\}$ defining $S$.
@@
    The extended projection SIC field $F^{\rm eproj}\!\left(S\right)=F^{\rm proj}\!\left(S\right)\!(\xi_{d})$ is obtained by adjoining the root of unity $\xi_{d}=-e^{\frac{\pi i}{d}}$.
@@
    For algebraic SICs (ones with $F^{\rm vec}\!\left(S\right)$ a finite extension of $\mathbb{Q}$), the Galois projection SIC field $F^{\rm Gproj}\!\left(S\right)$ is the normal closure of $F^{\rm eproj}\!\left(S\right)$.
We have the following field inclusions, with the last being valid only for algebraic SICs:
@@
The field $F^{\rm Gproj}\!\left(S\right)$ is an invariant of the Galois multiplet $[[S]]$ of an algebraic SIC $S$. In 4.8, we show that 1.1 implies the equalities
@@
The refinements of 1.1 formulated in this paper assert that Weyl–Heisenberg SICs in dimension $d\neq 3$ occur in Galois multiplets indexed by the intermediate orders $\mathcal{O}^{\prime}$ of $K_{d}$ that lie between $\mathcal{O}_{\Delta_{d}}=\mathbb{Z}[\varepsilon_{d}]$ and $\mathcal{O}_{K_{d}}$. Each intermediate order is in turn labeled by its conductor $f^{\prime}$, a positive integer such that $\operatorname{disc}(\mathcal{O}^{\prime})=(f^{\prime})^{2}\Delta_{0}$ where $\Delta_{0}=\operatorname{disc}(\mathcal{O}_{K_{d}})$. Quadratic orders are uniquely specified by their discriminants, so we may write $\mathcal{O}^{\prime}=\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}}$. The conductors appearing are the positive divisors $f^{\prime}$ of the conductor $f$ of $\mathcal{O}$.
@@
Fix a positive integer $d\neq 3$, let $\Delta=\Delta_{d}=(d+1)(d-3)$, and write $\Delta=f^{2}\Delta_{0}$ for $\Delta_{0}=\operatorname{disc}(\mathcal{O}_{K_{d}})$ and some positive integer $f$. Then there is a bijection $\mathcal{M}$ between intermediate orders $\mathcal{O}^{\prime}=\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}}$, labeled by their conductors $f^{\prime}$, and Galois multiplets,
@@
    The number of geometric equivalence classes $[S]$ in the multiplet $\mathcal{M}(f^{\prime})=[[S]]$ is the class number $h_{(f^{\prime})^{2}\Delta_{0}}:=\left|\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})\right|$ of the the order $\mathcal{O}^{\prime}$ of discriminant $(f^{\prime})^{2}\Delta_{0}$.
@@
    If $f_{1}\mid f_{2}$ and $f_{2}\mid f$, and $[S_{1}]\in\mathcal{M}(f_{1})$, and $[S_{2}]\in\mathcal{M}(f_{2})$, then $F^{\rm vec}\!\left(S_{1}\right)\subseteq F^{\rm vec}\!\left(S_{2}\right)$.
The special case $f^{\prime}=1$ corresponds to 1.1, where the order $\mathcal{O}^{\prime}$ is the maximal order. Appleby, Flammia, McConnell, and Yard [8] call the associated Galois multiplet $[[S]]$ the minimal multiplet, as it has the smallest associated field $F^{\rm vec}\!\left(S\right)$ under inclusion. The field inclusions asserted in 1.1(2) imply the truth of 1.2(2) in the special case $f_{1}=1$.
The following conjecture is an extension of both 1.1 and 1.2, which uses the notion of ray class fields of an order $\mathcal{O}^{\prime}$ introduced in [34] and reviewed in Section 6.1.
@@
Let $d\geq 4$. Then the map $\mathcal{M}$ in 1.2 may be chosen so that, for any Weyl–Heisenberg line-SIC $S$ in dimension $d$ with associated order $\mathcal{O}^{\prime}$ having discriminant $(f^{\prime})^{2}\Delta_{0}$ (that is, with $\mathcal{M}(f^{\prime})=[[S]]$), one has
$F^{\rm vec}\!\left(S\right)=F^{\rm proj}\!\left(S\right)=H_{d^{\prime}\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}^{\prime}},$ (1.6)
@@
The paper describes numerical checks of predictions of these conjectures against numerical data for SICs obtained in the datasets [41, 40, 22, 23]. The numerical data include both exact Weyl–Heisenberg SICs given by algebraic fiducial vectors and numerical SICs given as rational vectors approximately satisfying the fiducial vector SIC relations (to many decimal places). The notion of geometric equivalence class makes sense for numerical SICs, because they have well-defined $\operatorname{PEC}(d)$-orbits; however, the notion of Galois multiplet is not well-defined for them. In consequence we cannot test 1.2 using numerical SICs, since its statement involves Galois multiplets. Instead, we extract a numerical prediction implied by 1.2 that is testable.
This numerical prediction counts the total number of geometric equivalence classes of Weyl–Heisenberg SICs in dimension $d$ and is formulated as 5.4. It says that the total number of geometric equivalence classes of Weyl–Heisenberg SICs in dimension $d\geq 4$ is $\left|\operatorname{Clm}(\mathcal{O}_{\Delta_{d}})\right|$, the cardinality of the class monoid of the order $\mathcal{O}_{\Delta_{d}}$, as defined in Section 5.1.
We formulate this prediction as 5.4, and we test it in 5.7 using both exact and numerical SICs. These tests are summarized by Table 5.1 and Table 5.2 for all dimensions $4\leq d\leq 90$. 5.4 gives agreement in all these dimensions, assuming that the list of exact plus numerical SICs is complete. (The number $\left|\operatorname{Clm}(\mathcal{O}_{\Delta_{d}})\right|$ is equal to a sum of the ring class numbers $h_{(f^{\prime})^{2}\Delta_{0}}$ appearing in 1.2; see 5.3.)
The existence of the bijection $\mathcal{M}$ asserted in 1.2 implies a numerical prediction concerning the number of Galois multiplets, predicting that it equals the number of quadratic orders between $\mathcal{O}_{\Delta_{d}}$ and $\mathcal{O}_{K_{d}}$, which is the number of over-orders of $\mathcal{O}_{\Delta_{d}}$ in $K=\mathbb{Q}(\sqrt{\Delta_{d}})$. This numerical prediction is formulated as 5.8 and tested in 5.9 against certain dimensions $d$ in which all known SICs are exact SICs. Supporting data is given in Table 5.1.
In Section 6, we provide further tests of 1.2(2) and 1.3. We first summarize results on ray class fields of orders required to understand the conjectures. In 6.6, we numerically test the equality of degrees over $\mathbb{Q}$ of the ray class fields of orders in 1.3 to those of the fields $F^{\rm vec}\!\left(S\right)$. Tests are performed for exact SICs in dimensions $4\leq d\leq 15$ and $d=35$, and a supporting table of fields is given in Table 6.1.
We display for $d=35$ in 5.10 and 6.7 a (unique) bijective map $\mathcal{M}$ of orders to Galois multiplets that fulfills both conditions (1) and (2) of 1.2. In Section 8 we show that 1.3 strongly restricts the map $\mathcal{M}$ and, according to results proved elsewhere [36], determines it uniquely for a fraction $\frac{47}{48}$ of dimensions $d$. We show in 8.2 that the set of $d$ for which there exist two distinct orders $\mathcal{O}^{\prime},\mathcal{O}^{\prime\prime}$ with identical $H_{d^{\prime}\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}^{\prime}}$ in 1.3 is the set of $d$ for which the squarefree part of $(d+1)(d-3)$ is 1(mod $1\,\left(\operatorname{mod}\,8\right)$. Then 8.3 computes the asymptotic density of this set to be $\frac{1}{48}$. We present Table 8.1, which lists all cases of equality of such orders occurring for $d\leq 500$. These equality cases include all cases of observed equality of SIC fields $d\in\{47,67,259\}$ by Grassl [26].
@@ 1.4. Notation
  • Boldface roman lower case letters $\mathbf{v}=(v_{0},v_{1},\cdots v_{d-1})^{T}$ represent column vectors in $\mathbb{C}^{d}$. For a complex vector, $\mathbf{w}^{\dagger}:=\overline{\mathbf{w}}^{T}$, and similarly for matrices.
@@ 2. Weyl–Heisenberg SICs
In this section, we define and give background on Weyl–Heisenberg (covariant) SICs, which satisfy a particular group symmetry property.
@@ 2.1. Weyl–Heisenberg SICs
Specifically, Weyl–Heisenberg line-SICs will be defined as the orbit of a single line under the action of the Weyl–Heisenberg group, a finite subgroup of the group of $d\times d$ unitary matrices.
@@
The projective Weyl–Heisenberg group $\operatorname{PWH}(d)$ is defined to be $\operatorname{WH}(d)$ modulo its scalar matrix subgroup $Z(\operatorname{WH}(d))=\{\xi^{j}I:\,j\geq 0\}$, where $I$ is the $d\times d$ identity matrix, $X$ is called the cyclic shift matrix, and $Z$ is called the modulation matrix [46, Sec. 14.5].
@@
Its associated set of complex lines, denoted $\operatorname{WHL}(\mathbf{w})$, is
@@
We call a generating vector $\mathbf{v}$ of a Weyl–Heisenberg SIC a fiducial vector. Such a vector is well-defined up to multiplication by a complex scalar $z\in\mathbb{C}^{\times}$. Each SIC has a choice of $d^{2}$ fiducial vectors, up to scalars.
@@
This conjecture is currently known by explicit constructions to hold for finitely many dimensions $d$, including $1\leq d\leq 53$ and a extensive number of higher dimensions.
@@
For each dimension $d\geq 1$, except $d=3$, there are finitely many Weyl–Heisenberg SICs.
There are a few small dimensions where this conjecture has been proved rigorously, including $d\in\{1,2,4,5,6\}$. The cases $d=1,2$ may be done by hand; the cases $d=4,5,6$ have been handled by computer using Gröbner bases [27, unpublished]. In higher dimensions, up to at least $d=50$, it is supported by non-rigorous evidence given by numerical optimization routines repeatedly returning the same SICs [41, 40]; see Table 5.1. In dimension $3$, there is a one-parameter family of inequivalent Weyl–Heisenberg SICs.
A closely related conjecture is stated in Waldron [46, Sec. 14.27] as follows: The algebraic variety of SIC fiducials is zero dimensional (and nonempty), except for $d=3$, where it is one dimensional. Resolving 2.5 or Waldron’s conjecture in general (and specifically even for $n=7$) is an interesting open problem in optimization and real algebraic geometry.
@@ 2.2. Geometric equivalence classes of Weyl–Heisenberg SICs
There are two types of geometric operators acting on $\mathbb{C}^{d}$ that send SICs to SICs. Firstly, if $U$ is a unitary matrix in $\mathbf{U}(d)=\{U\in\mathbf{GL}(\mathbb{C}^{d}):UU^{\dagger}=1\}$, and $S=\{\mathbb{C}\mathbf{v}_{1},\ldots,\mathbb{C}\mathbf{v}_{d^{2}}\}$ is a SIC, then so is $US=\{\mathbb{C}U\mathbf{v}_{1},\ldots,\mathbb{C}U\mathbf{v}_{d^{2}}\}$. (This is a unitary equivalence.) Secondly, if $C_{d}$ is the complex conjugation operator acting pointwise on $\mathbb{C}^{d}$, so that $C_{d}\mathbf{v}:=\overline{\mathbf{v}}$, then $C_{d}$ also preserves the SIC property, as does any “antiunitary” operator of the form $C_{d}U$. These operators may be collected together to form the extended unitary group.
@@
The extended unitary group is $\mathbf{EU}(d):=\mathbf{U}(d)\sqcup C_{d}\mathbf{U}(d)$. The projective extended unitary group $\mathbf{PEU}(d):=\mathbf{EU}(d)/\{zI:z\in\mathbb{C},\left|z\right|=1\}$.
@@
If $\mathbf{v}$ is a fiducial vector for a Weyl–Heisenberg covariant SIC and $U\in\mathbf{EU}(d)$, then the vector $U\mathbf{v}$ will be a fiducial vector for a SIC covariant for the group $G^{\prime}=U\operatorname{WH}(d)U^{-1}$. All currently known SICs are unitary-equivalent to a Weyl–Heisenberg group $\operatorname{WH}(d)$ covariant SIC, with a single exception, the Hoggar lines in $\mathbb{C}^{8}$ [30, 31]. The Hoggar lines are group covariant for the tensor product $\operatorname{WH}(2)\otimes\operatorname{WH}(2)\otimes\operatorname{WH}(2)$. (The group $\operatorname{WH}(2)\otimes\operatorname{WH}(2)\otimes\operatorname{WH}(2)$ may itself be thought of as a generalized (Weyl)–Heisenberg group.)
@@
The Clifford group $\operatorname{C}(d)$ is infinite because it contains a copy of $\mathbf{U}(1)$ as scalar matrices $zI$ with $\left|z\right|=1$; thus, $\operatorname{EC}(d)$ is infinite. The groups $\operatorname{PC}(d)$ and $\operatorname{PEC}(d)$ are finite groups. (The extended Clifford $\operatorname{EC}(d)$ group is studied in connection with SICs by Appleby [1].)